# The graphene-water interface is acidic

### v1

@@ Abstract
Water’s ability to autoionize into hydroxide and hydronium ions profoundly influences surface properties, rendering interfaces either basic or acidic. While it is well-established that the water-air interface is acidic, a critical knowledge gap exists in technologically relevant surfaces like the graphene-water interface. Here we use machine learning-based simulations with first-principles accuracy to unravel the behavior of the hydroxide and hydronium ions at the graphene-water interface. Our findings reveal that the graphene-water interface is acidic, with the hydronium ion predominantly residing in the first contact layer of water. In contrast, the hydroxide ion exhibits a bimodal distribution, found both near the surface and towards the interior layers. Analysis of the underlying electronic structure reveals strong polarization effects, resulting in counterintuitive charge rearrangement. Proton propensity to the graphene-water interface challenges the interpretation of surface experiments and is expected to have far-reaching consequences for ion conductivity, interfacial reactivity, and proton-mediated processes.
@@
Water interacts with interfaces in numerous technologies involving areas such as atmospheric chemistry [1], water desalination [2], energy production via water splitting [3], and storage devices [4]. For these technologies, understanding the fundamental nature of these interfaces —whether they are acidic or basic— is essential to improve performance and facilitate scientific breakthroughs. At the molecular level, this is governed by the self-dissociation of water into hydroxide (OH-) and hydronium (H3O+) ions, which ultimately determines the pH of a solution and facilitates proton transfer. A deeper understanding of how surface interactions influence the propensity for hydronium and hydroxide ions would enable the optimization of these interfacial processes.
@@
Despite significant progress in understanding proton transfer in bulk water and microsolvation [5, 6, 7, 8, 9, 10, 11], describing these processes near interfaces continue to pose both theoretical and experimental challenges. One of the most enduring and fundamental debates in chemistry has been the acidity or basicity of the air-water interface [12, 13, 14, 15, 16]. The main complexity stems from the dynamic nature of protonic defects, namely hydronium and hydroxide ions, in the aqueous phase [17, 18, 19]. Furthermore, the interplay between directional hydrogen bonds and non-directional van der Waals forces often results in molecular conformations with similar energies [17, 18, 19], adding to the challenge. Experimental analysis is further complicated by varying probing resolutions and interfacial depths [14]. Only recently has it been established, both theoretically and experimentally, that the air-water interface is acidic, with the hydronium ion preferring the interface while the hydroxide ion does not [20, 21, 22, 23]. However, this understanding may not extend to other technologically relevant interfaces, as the fundamental mechanisms governing the behavior of these ions have been reported to differ across various interfaces [24, 25, 26, 27, 28, 29].
@@
Among all interfaces, the graphene-water interface is particularly relevant due to its extensive range of technological applications, from nanofluidics to electronic devices. In recent years, it has been reported that when water is confined between graphene sheets, it exhibits unique properties, including an anomalously low dielectric constant [30, 31], atypical friction behavior [32], and superionic character [33, 34], among other phenomena [35, 36]. Additionally, graphene’s precise synthesis in experimental settings [37] makes it an ideal candidate for providing insights applicable to real-world systems. However, our current understanding is limited, and previous studies have shown conflicting results. Grosjean et al. [38] reported a physisorbed state of hydroxide in the contact layer at the graphene/water interface, which is rationalized in the context of a macroscopic experimental observation[39] and has significant implications for conductance. Conversely, de Aquino et al. [40] indicated that hydroxide ions are more prevalent in the interior layers. Despite these efforts, a detailed picture of protonic defects near the graphene-water interface has not been established. In particular, whether the graphene-water interface is acidic or basic has remained unresolved until now. The complex interplay between water and the interface [41, 42], coupled with the prohibitive computational expense of ab initio molecular dynamics (AIMD) simulations needed to adequately sample these reactive systems, has significantly limited progress in this area. This study aims to address this gap, enhancing our understanding of graphene’s interfacial properties and improving the interpretation of experimental data.
@@
This work uses MLP-based simulations to demonstrate that the graphene-water interface is acidic, with hydronium ions predominantly residing at the interface, while hydroxide ions exhibit a bimodal distribution, being found both close to the surface and in layers farther from the interface. This surface affinity is due to the hydrogen bond environment of hydronium remaining stable at the interface, while hydroxide’s environment is disrupted. By examining the thermodynamic driving forces, we find that hydronium ions are enthalpically driven to the interface, whereas entropic forces drive hydroxide ions. When comparing these findings to the air-water interface, we see that graphene significantly influences ionic interactions due to polarization effects. This response suggests that macroscopic experiments should be interpreted carefully.
@@
To study protonic defects at the graphene-water interface, we developed an MLP using the MACE architecture [52]. The MLP demonstrates excellent capability in reproducing the potential energy surface of the underlying DFT at the revPBE-D3[53, 54] level of theory, known to perform well for water [55, 56, 57] and graphene-water[58] interactions, while also effectively capturing protonic defects [59]. Our model effectively captures the properties of both types of protonic defects in the water near free-standing graphene surfaces, under conditions ranging from ultra-confined environments to bulk-like settings (see Methods and Section S2 for development and validation details). We used the MLP to perform multi-nanosecond MD simulations, evaluating properties of five systems involving average slit widths of about 6.5, 9.2, 12.2, 14.7, 19.7 Å. These simulations included one layer (1L), two layers (2L), three layers (3L), four layers (4L), and five layers (5L) of water, each containing either a hydronium or a hydroxide ion (see Section S1).
@@
In Fig. 1(a), the water density profiles for the 1L–5L systems reveal distinct layers of water near graphene sheets, consistent with observations of water at solid surfaces [60, 61, 62, 63, 64]. Water forms sharply defined ‘interfacial’ layers in direct contact with the graphene sheets across all systems. For the thicker slits, we observe smoother ‘intermediate’ layers, and in the 5L system, these are accompanied by ‘bulk-like’ behavior as the bulk density of water is approached. To investigate the hydronium ion, we analyzed the density profiles from the specific oxygen of the protonic defect. As depicted in Fig. 1(b), the hydronium ion predominantly resides in the first contact layers of water, the interfacial water layers. Although the hydronium ion is still observed in the water layers further away from the interface, it appears less frequently there. For the hydroxide ion, shown in Fig. 1(c), the situation is less clear-cut: it can be present either at the interface or in the interior of the film, generally preferring the layers farther from the interface. These patterns are consistent across all studied slit widths featuring an intermediate water region (i.e., 3L–5L). Furthermore, the flexibility of graphene and its impact on these observations have been examined, confirming consistency even with completely rigid graphene sheets as shown in Section S3.
@@
Figure 2: Free energy profiles and water structuring of hydronium and hydroxide near graphene sheets. Free energy profiles for the hydronium and hydroxide ions as a function of their oxygen distance O to the closest graphene sheet obtained from the symmetrized density profiles. The minima in the interfacial layer, marked with a white dot, serve as the free energy reference point for each ion, with their distances to the interface presented. The structuring of the water layers is represented by the water density profiles, which are indicated with corresponding lighter colors and dashed lines. The shaded regions indicate the uncertainty obtained using the five replicate simulations. The background’s faded blue, orange, and gray represent the interfacial, intermediate, and bulk-like water layers, respectively. The horizontal axis is displayed up to half the distance between the two layers.
@@
To elucidate the molecular mechanism behind the observed ion behaviors, we investigate the structural characteristics of both defects, focusing on their hydrogen bonding patterns and orientational preferences at the interface. These factors are crucial for understanding their affinity for specific interfacial locations. In bulk water, the hydronium ion consistently donates three hydrogen bonds to neighboring water molecules while acting very rarely as an acceptor. Conversely, the hydroxide ion typically adopts a hypercoordinated square-planar arrangement in bulk water, accepting mostly four hydrogen bonds and transiently donating one. This arrangement, common in bulk environments, keeps the local water structure stable and enhances the ion’s stability [19]. In contrast to these well-established bulk solvation patterns, our analysis at the interface reveals significant differences. As shown in Fig. 3(a), the hydronium ion maintains a stable hydrogen bond environment across all layers, consistently donating three hydrogen bonds regardless of its location. Near the interface, the hydronium ion positions its hydrogen atoms toward the water layers, lying flat as shown in the snapshot. This orientation maintains the hydrogen bond network of the water molecules in the interior layers and is influenced by the hydrophobic nature of the hydronium ion’s oxygen, which typically does not accept hydrogen bonds due to its limited availability of lone-pair electrons. In contrast, as shown in Fig. 3(b), the hydroxide ion at the interface experiences significant hydrogen bond disruption, resulting in it accepting fewer hydrogen bonds than in bulk water and donating almost none due to spatial constraints near the interface. This disruption changes its typical orientation, with the hydroxide ion’s hydrogen atom predominantly facing the interface, impacting its usual fourfold hypercoordinated solvation pattern and significantly reducing its stability. Overall, this analysis shows that hydronium’s hydrogen bonding is not disrupted at the interface whereas the hydroxides’ is.
Figure 3: Difference in hydrogen bonding behavior of the protonic defects at the graphene-water interface. The average number of hydrogen bonds (accepted or donated, as indicated in each legend) for the hydronium ion (a) and the hydroxide ion (b) across different water layers, with representative snapshots showing their orientational preferences at the graphene-water interface. The error bars, smaller than the markers, are obtained using the five replicate simulations.
@@
Figure 4: Temperature dependence of the free energy for the protonic defects. (a) Free energy profiles for the hydronium and hydroxide ions as a function of their oxygen O distance to the closest graphene sheet at 300–400 K for the 3L system. The minima in the interfacial layers serve as the reference point for these profiles. The shaded regions represent the resulting uncertainty from five replicate simulations for each temperature and protonic defect, each propagated for 2.5 ns. (b) Adsorption free energies of the hydronium and hydroxide ions to the interface as a function of temperature, along with their respective linear fits. For the hydronium ion, we obtain $\Delta H=-2.4\pm 0.2$ kcal/mol and $\Delta S=3.3\pm 0.5$ cal/mol/K. For the hydroxide ion, we obtain $\Delta H=1.2\pm 0.2$ kcal/mol and $\Delta S=-3.0\pm 0.6$ cal/mol/K. The error bars, smaller than the markers, are obtained from the standard deviation of the five replicate simulations.
@@
The hydronium and hydroxide surface propensities reported herein bear similarities with those previously observed at the air-water interface, where the interface is typically described as being enriched with hydronium ions and depleted of hydroxide ions relative to the bulk. In our study, similar to what is observed at the air-water interface, hydronium ions are typically closer to the graphene layers than hydroxide ions, indicated by their shorter oxygen distances O from the sheet (recall Fig. 2). However, hydroxide ions at the graphene-water interface have a distinct local (but not global) free energy minimum at the interface, which is absent at the air-water interface. As discussed above, this difference in affinity arises from direct interactions between the liquid environment (including the protonic defects) and the graphene layers. Additionally, unlike at the air-water interface where ions can disrupt entropic surface stabilization through capillary wave actions, this effect is diminished or absent at the graphene-water interface, contributing to the differences between graphene-water compared to air-water [65].
@@
To assess the significance of the interactions between the liquid environment and the graphene layers, we used DFT to analyze the electron density of the interfaces (see definition in Methods). In particular, we looked at electron density differences, aimed at capturing the rearrangement of electron density due to the interaction between graphene and the liquid environment. Key results of this analysis are shown in Fig. 5 where it can be seen that graphene significantly alters the electron density of water molecules at the interface. This charge rearrangement —which is limited to the contact layers— therefore creates a distinct hydrogen bonding environment for protonic defects at the graphene-water interface compared to what they experience at the air-water interface. In addition, interesting and quite specific charge reorganization around the interfacial protonic defects themselves is observed. As shown in the insets of Fig. 5, the oxygen of the hydronium ion becomes less negatively charged upon interacting with graphene, and the nearby carbon atoms become slightly positively polarized by O (the latter effect is quantified by Bader charge analysis; see Section S5). When the hydroxide ion points with its hydrogen toward the interface, the graphene can cause a pronounced reorganization of the electron density, inducing electron accumulation over the entire C6 ring. Correspondingly, Bader charge analysis reveals that the hydroxide ion induces slight negative polarization on the nearby carbon atoms (see Section S5). This polarization is more subtle than that caused by the hydronium ion because the dangling hydrogen has more freedom in its orientation of the hydroxide ion leading to variable polarization effects. These findings demonstrate a clear outcome: the response of graphene to ions at the interface is counterintuitive. The cation induces charge depletion in nearby carbon atoms, while the anion induces charge accumulation. To understand this behavior is it crucial to consider their asymmetrical charge arrangements and preferred orientations at the interface. This intriguing finding emphasizes the need to carefully consider these intricate interactions when interpreting surface experiments at the graphene-water interface, such as zeta potential measurements.

### v2

# Protons accumulate at the graphene-water interface

@@ Abstract
Water’s ability to autoionize into hydroxide and hydronium ions profoundly influences surface properties, rendering interfaces either basic or acidic. While it is well-established that protons show an affinity to the air-water interface, a critical knowledge gap exists in technologically relevant surfaces like the graphene-water interface. Here we use machine learning-based simulations with first-principles accuracy to unravel the behavior of the hydroxide and hydronium ions at the graphene-water interface. Our findings reveal that protons accumulate at the graphene-water interface, with the hydronium ion predominantly residing in the first contact layer of water. In contrast, the hydroxide ion exhibits a bimodal distribution, found both near the surface and towards the interior layers. Analysis of the underlying electronic structure reveals local polarization effects, resulting in counterintuitive charge rearrangement. Proton propensity to the graphene-water interface challenges the interpretation of surface experiments and is expected to have far-reaching consequences for ion conductivity, interfacial reactivity, and proton-mediated processes.
@@
Water interacts with interfaces in numerous technologies involving areas such as atmospheric chemistry [1], water desalination [2], energy production via water splitting [3], and storage devices [4]. For these technologies, understanding the fundamental nature of these interfaces —whether they accumulate or repel ions — is essential to improve performance and facilitate scientific breakthroughs. At the molecular level, this is governed by the self-dissociation of water into hydroxide (OH-) and hydronium (H3O+) ions, which ultimately determines the pH of a solution and facilitates proton transfer. A deeper understanding of how surface interactions influence the propensity for hydronium and hydroxide ions would enable the optimization of these interfacial processes.
@@
Despite significant progress in understanding proton transfer in bulk water and microsolvation [5, 6, 7, 8, 9, 10, 11], describing these processes near interfaces continue to pose both theoretical and experimental challenges. One of the most enduring and fundamental debates in chemistry has been the nature of the excess proton (hydronium ion) and hydroxide ion at the air-water interface [12, 13, 14, 15, 16]. The main complexity stems from the dynamic nature of protonic defects, namely hydronium and hydroxide ions, in the aqueous phase [17, 18, 19]. Furthermore, the interplay between directional hydrogen bonds and non-directional van der Waals forces often results in molecular conformations with similar energies [17, 18, 19], adding to the challenge. Experimental analysis is further complicated by varying probing resolutions and interfacial depths [14]. Only recently has it been established, both theoretically and experimentally, that the air-water interface accumulates protons, while hydroxide ions are repelled [20, 21, 22, 23]. However, this understanding may not extend to other technologically relevant interfaces, as the fundamental mechanisms governing the behavior of these ions have been reported to differ across various interfaces [24, 25, 26, 27, 28, 29].
@@
Among all interfaces, the graphene-water interface is particularly relevant due to its extensive range of technological applications, from nanofluidics to electronic devices. In recent years, it has been reported that when water is confined between graphene sheets, it exhibits unique properties, including an anomalously low dielectric constant [30, 31], atypical friction behavior [32], and superionic character [33, 34], among other phenomena [35, 36]. Additionally, graphene’s precise synthesis in experimental settings [37] makes it an ideal candidate for providing insights applicable to real-world systems. However, our current understanding is limited, and previous studies have shown conflicting results. Grosjean et al. [38] reported a physisorbed state of hydroxide in the contact layer at the graphene/water interface, which is rationalized in the context of a macroscopic experimental observation[39] and has significant implications for conductance. Conversely, de Aquino et al. [40] indicated that hydroxide ions are more prevalent in the interior layers, while hydronium ions are more prevalent in the interfacial layers. More recently, Scalfi et al. [41] revealed that hydronium adsorbs at the graphene-water interface while hydroxide mostly shows only limited adsorption under specific conditions. While these studies have advanced our understanding, they rely primarily on non-reactive force fields, which may constrain the mechanistic insights they can offer. As a result, despite valuable efforts in the field, a detailed picture of protonic defects near the graphene-water interface has not been established. In particular, detailed mechanistic and thermodynamic insights into hydrogen bonding, the orientational behavior, and interfacial polarization of the hydronium and hydroxide ion at the graphene-water interface have remained unresolved until now. The complex interplay between water and the interface [42, 43], coupled with the prohibitive computational expense of ab initio molecular dynamics (AIMD) simulations needed to adequately sample these reactive systems, has significantly limited progress in this area. This study aims to address this gap, enhancing our understanding of graphene’s interfacial properties and improving the interpretation of experimental data.
@@
This work uses MLP-based simulations to demonstrate that protons accumulate at the graphene-water interface, while hydroxide ions exhibit a bimodal distribution, being found both close to the surface and in layers farther from the interface. This surface affinity is due to the hydrogen bond environment of hydronium remaining stable at the interface, while hydroxide’s environment is disrupted. By examining the thermodynamic driving forces, we find that hydronium ions are enthalpically driven to the interface, whereas entropic forces drive hydroxide ions. When comparing these findings to the air-water interface, we see that graphene significantly influences ionic interactions due to polarization effects. This response suggests that macroscopic experiments should be interpreted carefully.
@@
To study protonic defects at the graphene-water interface, we developed an MLP using the MACE architecture [53]. The MLP demonstrates excellent capability in reproducing the potential energy surface of the underlying DFT at the revPBE-D3[54, 55] level of theory, known to perform well for water [56, 57, 58] and graphene-water[59] interactions, while also effectively capturing protonic defects [60]. We have verified the validity of our results with respect to the functional and dispersion correction and confirmed that the observed trends remain unchanged (see Section S2). Our model effectively captures the properties of both types of protonic defects in the water near free-standing graphene surfaces, under conditions ranging from ultra-confined environments to bulk-like settings (see Methods and Section S2 for development and validation details). We used the MLP to perform multi-nanosecond MD simulations, evaluating properties of five systems involving average slit widths of about 6.5, 9.2, 12.2, 14.7, 19.7 Å. These simulations included one layer (1L), two layers (2L), three layers (3L), four layers (4L), and five layers (5L) of water, each containing either a hydronium or a hydroxide ion (see Section S1).
@@
In Fig. 1(a), the water density profiles for the 1L–5L systems reveal distinct layers of water near graphene sheets, consistent with observations of water at solid surfaces [61, 62, 63, 64, 65]. Water forms sharply defined ‘interfacial’ layers in direct contact with the graphene sheets across all systems. For the thicker slits, we observe smoother ‘intermediate’ layers, and in the 5L system, these are accompanied by ‘bulk-like’ behavior as the bulk density of water is approached (see Section S2). To investigate the hydronium ion, we analyzed the density profiles from the specific oxygen of the protonic defect. As depicted in Fig. 1(b), the hydronium ion predominantly resides in the first contact layers of water, the interfacial water layers. Although the hydronium ion is still observed in the water layers further away from the interface, it appears less frequently there. For the hydroxide ion, shown in Fig. 1(c), the situation is less clear-cut: it can be present either at the interface or in the interior of the film, generally preferring the layers farther from the interface. These patterns are consistent across all studied slit widths featuring an intermediate water region (i.e., 3L–5L). Furthermore, the flexibility of graphene and its impact on these observations have been examined, confirming consistency even with completely rigid graphene sheets as shown in Section S3.
@@
Figure 2: Free energy profiles and water structuring of hydronium and hydroxide near graphene sheets. Free energy profiles for the hydronium and hydroxide ions as a function of their oxygen distance O to the closest graphene sheet obtained from the symmetrized density profiles. The minima in the interfacial layer, marked with a white dot, serve as the free energy reference point for each ion, with their distances to the interface presented. The structuring of the water layers is represented by the water density profiles, which are indicated with corresponding lighter colors and dashed lines. The shaded regions indicate the uncertainty resulting from the standard deviation of the five replicate simulations. The background’s faded blue, orange, and gray represent the interfacial, intermediate, and bulk-like water layers, respectively. The horizontal axis is displayed up to half the distance between the two layers.
@@
To elucidate the molecular mechanism behind the observed ion behaviors, we investigate the structural characteristics of both defects, focusing on their hydrogen bonding patterns and orientational preferences at the interface. These factors are crucial for understanding their affinity for specific interfacial locations. In bulk water, the hydronium ion consistently donates three hydrogen bonds to neighboring water molecules while acting very rarely as an acceptor. Conversely, the hydroxide ion typically adopts a hypercoordinated square-planar arrangement in bulk water, accepting mostly four hydrogen bonds and transiently donating one. This arrangement, common in bulk environments, keeps the local water structure stable and enhances the ion’s stability [19]. In contrast to these well-established bulk solvation patterns, our analysis at the interface reveals significant differences. As shown in Fig. 3(a), the hydronium ion maintains a stable hydrogen bond environment across all layers, consistently donating three hydrogen bonds regardless of its location. Near the interface, the hydronium ion positions its hydrogen atoms toward the water layers, lying flat as shown in the snapshot. This orientation maintains the hydrogen bond network of the water molecules in the interior layers (see Section S5) and is influenced by the hydrophobic nature of the hydronium ion’s oxygen, which typically does not accept hydrogen bonds due to its limited availability of lone-pair electrons. In contrast, as shown in Fig. 3(b), the hydroxide ion at the interface experiences significant hydrogen bond disruption, resulting in it accepting fewer hydrogen bonds than in bulk water and donating almost none due to spatial constraints near the interface. This disruption changes its typical orientation, with the hydroxide ion’s hydrogen atom predominantly facing the interface, impacting its usual fourfold hypercoordinated solvation pattern and significantly reducing its stability. Overall, this analysis shows that hydronium’s hydrogen bonding is not disrupted at the interface whereas the hydroxides’ is.
Figure 3: Difference in hydrogen bonding behavior of the protonic defects at the graphene-water interface. Average number of hydrogen bonds (accepted or donated, as indicated in each legend) for the hydronium ion (a) and the hydroxide ion (b) across the different water layers with their corresponding bulk reference indicated with horizontal lines. The representative snapshots show their orientational preferences at the graphene-water interface. Hydrogen bonds are counted using the geometric definition provided in Ref. 66. The error bars, smaller than the markers, are obtained from the standard deviation of the five replicate simulations.
@@
Figure 4: Temperature dependence of the free energy for the protonic defects. (a) Free energy profiles for the hydronium and hydroxide ions as a function of their oxygen O distance to the closest graphene sheet at 300–400 K for the 3L system. The minima in the interfacial layers serve as the reference point for these profiles. The shaded regions represent the resulting uncertainty obtained from the standard deviation of the five replicate simulations for each temperature and protonic defect, each propagated for 2.5 ns. (b) Adsorption free energies of the hydronium and hydroxide ions to the interface as a function of temperature, along with their respective linear fits. For the hydronium ion, we obtain $\Delta H=-2.4\pm 0.2$ kcal/mol and $\Delta S=-3.0\pm 0.6$ cal/mol/K. For the hydroxide ion, we obtain $\Delta H=1.2\pm 0.3$ kcal/mol and $\Delta S=3.1\pm 0.8$ cal/mol/K. The error bars, smaller than the markers, are obtained from the standard deviation of the five replicate simulations.
@@
The hydronium and hydroxide surface propensities reported herein bear similarities with those previously observed at the air-water interface[21, 67], where the interface is typically described as being enriched with hydronium ions and depleted of hydroxide ions relative to the bulk. In our study, similar to what is observed at the air-water interface, hydronium ions are typically closer to the graphene layers than hydroxide ions, indicated by their shorter oxygen distances O from the sheet (recall Fig. 2). However, hydroxide ions at the graphene-water interface have a distinct local (but not global) free energy minimum at the interface, which is absent at the air-water interface. As discussed above, this difference in affinity arises from direct interactions between the liquid environment (including the protonic defects) and the graphene layers. Additionally, unlike at the air-water interface where ions can disrupt entropic surface stabilization through capillary wave actions, this effect is diminished or absent at the graphene-water interface, contributing to the differences between graphene-water compared to air-water [68].
@@
To assess the significance of the interactions between the liquid environment and the graphene layers, we used DFT to analyze the electron density of the interfaces (see definition in Methods). In particular, we looked at electron density differences, aimed at capturing the rearrangement of electron density due to the interaction between graphene and the liquid environment. Key results of this analysis are shown in Fig. 5 where it can be seen that graphene significantly alters the electron density of water molecules at the interface. This charge rearrangement —which is limited to the contact layers— therefore creates a distinct hydrogen bonding environment for protonic defects at the graphene-water interface compared to what they experience at the air-water interface. In addition, interesting local charge reorganization around the interfacial protonic defects is observed. As shown in the insets of Fig. 5, the oxygen of the hydronium ion exhibits a localized decrease in its negative charge upon interacting with graphene, while the nearby carbon atoms near O show a slight positive polarization. This localized effect is further quantified through Bader charge analysis (see Section S6). When the hydroxide ion points with its hydrogen toward the interface, it can induce electron accumulation above the entire C6 ring. However, Bader charge analysis reveals that this is only a local effect rather than a global charge reorganization (see Section S6). This polarization is more subtle than that caused by the hydronium ion because the dangling hydrogen has more freedom in its orientation of the hydroxide ion leading to variable polarization effects. Notably, while water molecules may occasionally orient their OH bond toward the graphene and induce similar polarization effects, this alignment is transient. In contrast, hydroxide ions consistently orient their OH group toward the graphene, resulting in a stronger and more persistent impact due to their stable interaction with the surface (see Section S6).
These findings demonstrate a clear outcome: the response of graphene to ions at the interface is counterintuitive. The cation induces charge depletion in nearby carbon atoms, while the anion induces charge accumulation. To understand this behavior is it crucial to consider their asymmetrical charge arrangements and preferred orientations at the interface. This intriguing finding emphasizes the need to carefully consider these intricate interactions when interpreting surface experiments at the graphene-water interface, such as zeta potential measurements.